puunlued annually for fi better? i) Compound Interest $6,766.45; Simple interest $6,537.50 Simple interest $7,124.32 c)Supposing that you are obtaining a personal loan from your uncle i be repaid in three years to cover some of your college expenses. , calculate the amount he would pay in three years time. 4. Solve the following equations 1, F = P(1 + i) P = $10,000 i = 4.5% n=6years equation 1 2. P= equation 2 (1 + F $30,000 i-1090 n=3years a. P=$7,000 i = 5% nl 3 years n2= 1 year b.) Briefly discuss the four 5a) As real estate investors, p fundamental principles of Engineering Eco ors, Max and Alona invested $90,000 in develc ssuming an interest rate of 4%, calculate the pre Alona in
Answer #1

4.

1. F=P(1+i)ⁿ

Or, F = 10000(1+0.045)^6 = $13,022.60125

2. \large P=\frac{F} {(1+i)^n}

So, \large P=\frac{30000} {(1+0.10)^3} = $22,539.444

3. F=P(1+i)^-n

So, F1 = 7000 * (1+0.05)^-3 = $6,046.8632

And,F2 = 7000 * (1+0.05)^-1 = $6,666.667

F = F1+F2 = $12,713.53